40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 18 04:04:27 2011, kiwamu, Update, LSC, measurement of sensing matrix : touchy SRM 
    Reply  Wed Oct 19 05:15:44 2011, kiwamu, Update, SUS, taking care of SRM f2pSRM.png
       Reply  Wed Oct 19 15:46:49 2011, kiwamu, Update, SUS, taking care of SRM SRM.png
Message ID: 5699     Entry time: Wed Oct 19 15:46:49 2011     In reply to: 5691
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: taking care of SRM 

Quote from #5691

I am going to check the sensing matrix with the new free swinging spectra (#5690)

 

The SRM input matrix has been readjusted.

However still there is the unwanted coupling from the POS drive to SIDE signal and from the SIDE drive to POS signal.

      BADNESS
  SRM  SRM.png       pit     yaw     pos     side    butt
UL    0.871   0.975   1.115  -0.295   1.096  
UR    1.015  -1.025   1.128  -0.140  -1.053  
LR   -0.985  -0.984   0.885  -0.088   0.831  
LL   -1.129   1.016   0.872  -0.243  -1.020  
SD    0.084   0.061   3.534   1.000  -0.018  
 
 4.24965

 

ELOG V3.1.3-